# Generative Verifiers: Reward Modeling as Next-Token Prediction

### v1

@@ 5 Related Work
Reward models (RMs) and verifiers. Conventionally, RMs and verifiers are trained as discriminative models via binary classification: given a prompt and a corresponding solution or a pair of solutions), the model is either trained to predict the correctness of the solution (Cobbe et al., 2021; Saunders et al., 2022; Lightman et al., 2023; Wang et al., 2023; Uesato et al., 2022; Luo et al., 2024; Yu et al., 2024) or a preference between the two solutions (Stiennon et al., 2020; Nakano et al., 2021). Concretely, the RM or verifier directly produces a numerical continuous-valued score, which is then plugged into a classification objective (3). As such, discriminative verifiers do not utilize the text generation capabilities of LLMs. In contrast to discriminative RMs, GenRM does not train RMs that output a numerical score from a special logit, but rather represent the correctness decision using the log probability of ‘Yes’ and ‘No’ tokens under special instructions. Posing verification as generating “yet another token” allows it to tap better into the generation capabilities of LLMs, by making it straightforward to employ CoT reasoning and additional inference-time compute for better verification.
@@
LLM-as-a-Judge for verification. Another line of work that poses verification as next-token prediction simply prompts off-the-shelf LLMs to act as a verifier when provided with a rubric and a template for grading (Zheng et al., 2024; Bai et al., 2022; Kim et al., 2023; Ling et al., 2024), but without any specific training for the same. Perhaps unsurprisingly, we find in our experiments that using substantially more powerful LLMs (Gemini 1.0 Pro) as a judge is substantially worse than our trained GenRM (using weaker Gemma models), highlighting the necessity of training generative verifiers. More generally, even the strongest proprietary LLMs to date, such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2024b), fall behind trained RMs on popular leaderboards, such as RewardBench (Lambert et al., 2024), and this gap is much larger for reasoning problems. While Agarwal et al. (2024) utilized many-shot prompting on an off-the-shelf LLM to obtain a “Yes/No” verifier, our work focuses on training generative verifiers and also leverages CoT rationales and test-time compute.
@@
Using CoTs for reward models. Prior works have also considered using critiques or CoT to extract preference and verification signals (Yuan et al., 2024; Wu et al., 2024; Wang et al., 2024; Ye et al., 2024a); in contrast to these prior works, GenRM utilizes model-generated CoT directly for training the verifier. Upon inference, a GenRM produces its own CoTs, which it then uses to make decisions on correctness, unlike other work that simply uses CoTs from a separate (and highly-capable) model (Ye et al., 2024b). Compared to (McAleese et al., 2024) which assumes access to high-quality data from humans to train discriminative RMs for generating code critiques, we show that GenRM can be trained from purely synthetic, model-generated critiques.
Concurrent work (Ankner et al., 2024) trains a model to produce response critiques, which are then passed as input into a reward modelling head, separate from the base LLM. Unlike GenRM which uses next-token prediction, their RM head is trained discriminatively akin to standard RMs. While this approach allows them to leverage CoTs and inference-time computation, it does not allow them to unify solution generation and verification together as a result of a discriminative RM head, that GenRM seamlessly enables (Section 4.2). Moreover, our results show that even using GenRM to produce a single verification token outperforms standard RMs even without CoT or inference-time compute, which can potentially improve the results in this concurrent work (Figure 13, Left).
@@
Unified generation and verification. One of the hallmark properties of GenRM is that the very same generative verifier can be co-trained with a generation objective (5): when given a problem, the model is trained to produce a solution, whereas when given a problem and a candidate solution, it is trained to verify this candidate. This is related to DPO (Rafailov et al., 2024) and its application to learning verifiers in reasoning (Hosseini et al., 2024), which aims to unify generation (policy) and verification (reward models) by representing the reward implicitly using the logits of a policy and training the policy with a reward-modelling loss. For reasoning, this type of model tying has been shown to exhibit erroneous extrapolation and a degradation in learned representations, which prior work has attempted to address with additional techniques (Pang et al., 2024; Setlur et al., 2024; Pal et al., 2024; Yang et al., 2024). Of these, while the approach of Yang et al. (2024) trains an reward model with an auxiliary generative SFT loss, note that this loss is applied on a separate head for regularization purposes and is discarded after training; unlike GenRM no text is produced when querying the RM.
Unlike DPO, our unification approach is distinct: GenRM treats verification as next-token prediction, allowing the same model to be trained on both verification and policy tasks using their respective prompts. This eliminates the need for separate reward modeling or loss modifications, offering greater flexibility for practitioners to combine various reward and generation objectives. For example, while Hosseini et al. (2024) struggled to obtain a DPO verifier that can generate correct solutions, the GenRM-CoT verifier has better generation performance compared to SFT only on correct solutions (Figure 7).
@@ 6 Conclusion & Future Work
In this paper, we have introduced Generative Verifiers (GenRM), which recast verification as next-token prediction in LLM reasoning domains. GenRM is a more performant alternative to discriminative reward models, and unlocks the use of powerful tools, such as chain-of-thought reasoning and majority voting for better verification. GenRM also unifies generation and verification into a single LLM, and demonstrates that such a unification benefits both generation and verification. Moreover, we show that synthetic model-generated rationales, which are noisy and sub-optimal in many aspects, are sufficient to teach GenRM how to use verification CoT to pick out tricky reasoning errors in grade school math problems. Future work includes extending the generative verification framework to a broader range of tasks (such as coding, alignment, and answering open-ended questions), and studying how generative verifiers can be integrated into existing LLM self-improvement algorithms (Singh et al., 2023; Gulcehre et al., 2023).
As a generative LLM for verification, GenRM offers a solid foundation for future work. It’s readily compatible with all the existing tools designed to improve LLMs, such as retrieval-augmented generation (Borgeaud et al., 2022), many-shot learning (Agarwal et al., 2024), multi-staged prompting (Yao et al., 2024), tool use (Schick et al., 2024), and code generation and execution (Chen et al., 2021). Future work can explore how those techniques can be in service of verifying model-generated outputs under the generative verifier framework.
@@ Acknowledgements
This work was done during LZ, AH, and HB’s internship at Google. We thank Hugo Larochelle, Minqi Jiang, Aleksandra Faust, Ankit Anand, Guillaume Desjardins, Doina Precup, and Charlie Snell for feedback on an earlier version of this paper and informative discussions. We thank Chirag Nagpal and Katrin Tomanek for support in setting up infrastructure for experimentation that was crucial for running our experiments.
@@ Appendix A Training Data Generation for Verifiers
  • Last Letter Concatenation: To generate the training data, for each length $\{2,3,4\}$, we generate $350$ problem queries by randomly sampling from the set of words in original training set; for each problem query, we generate 128 attempts from Gemma-2B (Team et al., 2024a) model. This gives us a total of about 50K training data points after de-duplication. We train verifiers on examples of lengths $\{2,3,4\}$ (here the length refers to how many words are in the input list), and evaluate the verifier performance on length 6. We use the format in Table A.1 to algorithmically generate ground-truth verification CoT for training.
  • Word Sorting: We train verifiers on a dataset comprised of $\{2,3,4\}$ words in each example, and evaluate the performance on length $5$. For each length, we generate 4096 lists of words as the problem queries; for each problem, we generate 64 attempts from Gemma-2B. After de-duplication and filtering out invalid responses, we have a total of about 100K training data points. We also algorithmically generate ground-truth verification CoT for training (see Table A.1).
  • Grade School Math: We follow the original train/test split and use 1.3K problems for test, 128 problems for validation, and about 7.2K problems for training. We generate 50 solutions per problem, and randomly sample at max 16 correct solutions and 16 incorrect solutions per problem as the training set.
@@
Table A.2: We use model-generated rationales as CoT training data on GSM with the above prompt. When generating the training data for rationales, we show the model another model-generated solution that arrives at the correct answer, which is privileged information that does not exist at test time. Our method does not require a more capable model for the rationales compared to the solutions: we use the same model to generate solutions and synthetic rationales in the training data.
@@ Appendix B Hyper-parameters for Verifier Training
We finetune Gemma-based language models. After doing a sweep of learning rates (LR), we find that an LR of $[2e-6,1e-6,5e-7]$ works well for our tasks considered (with LR=$2e-6$ generally being the best). We use a weight decay of $21e-$ 2, and do not apply any dropout. We use the Adam optimizer (Kingma, 2014) with decoupled weight decay (Loshchilov and Hutter, 2017) and a gradient norm clipping of $1.0$. We use a linear warmup of $1000$ gradient steps, and a cosine decay schedule that decays to $10\%$ of the peak learning rate towards the end of training. We finetune for 200K steps with a batch size of 128, and use seqio (Roberts et al., 2022) library to create data mixtures. We pick the best checkpoint based on validation accuracy of verification on held out problems and solutions. We always use data balancing between 50% correct solutions and 50% incorrect solutions in training.
@@
We finetune Gemma-based discriminative RMs by using a special token’s logit for classification. We chose the best performing ORM on our validation sets by launching a large sweep over learning rates $[1e-7,5e-7,1e-6,2e-6,3e-6,5e-6]$, weight decay $[1e-3,1e-2,1e-1]$ and dropouts $[1e-3,1e-2,4e-2]$. We also schedule the learning rate with a linear ramp up and a cosine decay. We then pick the best model according to validation accuracy of verification on held out problems and solutions. We always use data balancing between 50% correct solutions and 50% incorrect solutions in training.
Appendix C Examples of verification rationales generated by GenRM-CoT on GSM8K
Example of on Grade School Math
@@ Appendix C Examples of verification rationales generated by GenRM-CoT on GSM8K
Table C.1: GenRM CoT Example 1
@@
Table C.2: GenRM CoT Example 2
@@
Table C.3: GenRM CoT Example 3
@@
Table C.4: GenRM CoT Example 3 (Continued)
@@
Table C.5: GenRM CoT Example 4
@@
Table C.6: GenRM CoT Example 5
@@
Table C.7: GenRM CoT Example 6
@@
Table C.8: GenRM CoT Example 7
@@
Table C.9: GenRM CoT Example 8
@@
Table C.10: GenRM CoT Example 9
Example of GenRM-CoT on Grade School Math
Discriminative RM score: 0.8414
GenRM-CoT: Let’s verify step by step.
**Verification:**
Step 1: Half of 4 hours is 2 hours.
Expected Answer: Correct.
Verification: Is the answer correct (Yes/No)? No
Gen-RM (Majority Voting) score: 0.2498
Table C.11: GenRM CoT Example 10
@@
Table C.12: GenRM CoT Example 11

### v2

@@ 5 Related Work
Reward models (RMs) and verifiers. Conventionally, RMs and verifiers are trained as discriminative models via binary classification: given a prompt and a corresponding solution or a pair of solutions), the model is either trained to predict the correctness of the solution (Cobbe et al., 2021; Lightman et al., 2023; Wang et al., 2023; Uesato et al., 2022; Luo et al., 2024; Yu et al., 2024) or a preference between the two solutions (Stiennon et al., 2020; Nakano et al., 2021). Concretely, the RM directly produces a numerical continuous-valued score, which is then plugged into a classification objective (3). As such, discriminative verifiers do not utilize the generation capabilities of LLMs. In contrast to discriminative RMs, GenRM represents the correctness decision using the log probability of specific tokens, for example ‘Yes’ and ‘No’. Posing verification as generating “yet another token” allows it to tap better into the generation capabilities of LLMs, by making it straightforward to employ CoT reasoning and additional inference-time compute for better verification.
@@
LLM-as-a-Judge. Another line of work that poses verification as next-token prediction simply prompts off-the-shelf LLMs to act as a verifier when provided with a rubric and a template for grading (Zheng et al., 2024; Bai et al., 2022; Kim et al., 2023; Ling et al., 2024) or many-shot ICL examples (Agarwal et al., 2024), but without any specific training for the same. Perhaps unsurprisingly, we find in our experiments that using more powerful LLMs (Gemini 1.0 Pro) as a judge is worse than our trained GenRM using weaker Gemma models (Figure 1, 11), highlighting the necessity of training generative verifiers. Our generative verifiers also exhibit good out-of-distribution generalization, which might be due to better calibrated uncertainty estimates from training (Kapoor et al., 2024). More generally, even the strong proprietary LLMs, such as GPT-4 (Achiam et al., 2023) and Gemini  (Team et al., 2024), fall behind trained RMs on popular leaderboards (Lambert et al., 2024), and this gap is much larger for reasoning.
@@
Using CoTs for reward models. Prior works have also used critiques or CoT to extract preference and verification signals using LLM-as-a-Judge (Yuan et al., 2024; Wu et al., 2024; Wang et al., 2024); in contrast to these works, GenRM utilizes model-generated CoTs directly for training the verifier. Upon inference, a GenRM-CoT produces its own CoTs, which it then uses to make decisions on correctness, unlike Ye et al. (2024) that simply uses CoTs from a separate highly-capable LLM. In contrast to prior work that utilizes high-quality data from humans to train critique models (Saunders et al., 2022) or train discriminative RMs for generating code critiques (McAleese et al., 2024), we show that GenRM can be trained from purely synthetic, model-generated critiques. Concurrent work (Ankner et al., 2024) trains an RM to produce response critiques for preference pairs generated using a much more capable LLM, which are then passed as input into a RM head, separate from the base LLM. Unlike GenRM which uses next-token prediction, their RM head is trained discriminatively akin to standard RMs. While this approach allows them to leverage CoT, it does not allow them to unify solution generation and verification as a result of a discriminative RM head, which GenRM seamlessly enables (Section 4.2). Moreover, their synthetic critiques are not filtered for correctness, which would lead to poor verification CoTs on reasoning tasks (§3.3).
@@
Unified generation and verification. One of the hallmark properties of GenRM is that the same generative verifier can be co-trained with a generation objective (5): when given a problem, the model is trained to produce a solution, whereas when given a problem and a candidate solution, it is trained to verify this candidate. This is related to DPO (Rafailov et al., 2024) and its application to learning verifiers in reasoning (Hosseini et al., 2024), which aims to unify generation (policy) and verification (reward models) by representing the reward implicitly using the logits of a policy and training the policy with a reward-modeling loss. For reasoning, this type of model tying has been shown to exhibit erroneous extrapolation and degradation in learned representations, which prior work has attempted to address with additional techniques (Pang et al., 2024; Setlur et al., 2024; Pal et al., 2024; Yang et al., 2024). Of these, while Yang et al. (2024) train a reward model with an auxiliary generative SFT loss, note that this loss is applied on a separate head for regularization purposes and is discarded after training; unlike GenRM no text is produced when querying the RM. In addition, compared to DPO, GenRM uses a simpler next-token prediction loss, does not require a reference policy, and obtains significantly better verification performance (Figure 1, 7).
@@ 6 Conclusion & Future Work
In this paper, we have introduced Generative Verifiers (GenRM), which recast verification as next-token prediction. GenRM is more performant than discriminative verifiers, and unlocks the use of chain-of-thought reasoning and inference-time compute for better verification. GenRM also unifies generation and verification into a single LLM, and demonstrates that such a unification benefits both generation and verification. Moreover, we show that synthetic model-generated rationales, which can be error-prone, are sufficient to teach GenRM how to use verification CoT to pick out tricky errors on math reasoning tasks (see Figure 2, 4, 14, and Appendix E).
The framework of generative verification offers a solid foundation for future work. Promising directions include extending this framework to broader tasks such as coding, alignment, text-to-image generation (Lin et al., 2024), and open-ended generation (Besta et al., 2024). Furthermore, leveraging process-level supervision (Lightman et al., 2023) and training CoT verifiers with reinforcement learning (RL) can result in more accurate generative verifiers. Given GenRM’s compatibility with all the existing tools designed to improve LLMs, exploring enhancements through techniques like retrieval-augmented generation (Borgeaud et al., 2022), many-shot learning (Agarwal et al., 2024), multi-staged prompting (Yao et al., 2024), and tool use (Schick et al., 2024) would be interesting. Finally, incorporating generative verifiers into RL pipelines for LLMs warrants further investigation.
@@ Acknowledgements
This work was done during LZ, AH, and HB’s internship at Google. We thank Hugo Larochelle, Minqi Jiang, Aleksandra Faust, Ankit Anand, Guillaume Desjardins, Doina Precup, and Charlie Snell for feedback on an earlier version of this paper and informative discussions. We thank Chirag Nagpal and Katrin Tomanek for support in setting up infrastructure that was crucial for running Gemma experiments.
Author Contributions
LZ led the project, and ran almost all of the experiments and ablation studies, and wrote and edited the paper. AH was responsible for the discriminative RM baselines and DPO baselines. HB was responsible for the word-sorting task, and helped set up evaluations and DPO. MK advised the project, and provided feedback on writing. AK conceived the project with RA, advised LZ, and provided feedback on paper. RA hosted LZ as a student researcher, proposed several ideas and experiments, implemented reference-guided grading, wrote the initial draft and advised the project.
@@ Appendix A Training Data Generation for Verifiers
  • Last Letter Concatenation (Wei et al., 2022): Given a list of words, the task is to concatenate the last letters of each word (for instance, “Noah Paul Elisha Rebecca” $\rightarrow$ “hlaa”). To generate the training data, for each length $\{2,3,4\}$, we generate $350$ problem queries by randomly sampling from the set of words in original training set; for each problem query, we generate 128 attempts from Gemma-2B (Gemma Team et al., 2024a) model. This gives us a total of about 50K training data points after de-duplication. We train verifiers on examples of lengths $\{2,3,4\}$ (here the length refers to how many words are in the input list), and evaluate the verifier performance on length 6. We use the format in Table A.1 to algorithmically generate ground-truth verification CoT for training.
  • Word Sorting (Suzgun et al., 2022): Given a list of words, sort them in alphabetical order. We train verifiers on a dataset comprised of $\{2,3,4\}$ words in each example, and evaluate the performance on length $5$. For each length, we generate 4096 lists of words as the problem queries; for each problem, we generate 64 attempts from Gemma-2B. After de-duplication and filtering out invalid responses, we have a total of about 100K training data points. We also algorithmically generate ground-truth verification CoT for training (see Table A.1).
  • Grade School Math (Cobbe et al., 2021): We follow the original train/test split and use 1.3K problems for test, 128 problems for validation, and about 7.2K problems for training. We generate 50 solutions per problem, and randomly sample at max 16 correct solutions and 16 incorrect solutions per problem as the training set. We evaluate the verifier performance on 16 solutions per problem in the test set.
Table A.2: We use model-generated rationales as CoT training data on GSM with the above prompt with Gemini 1.0 Pro. Specifically, we show the model another solution that arrives at the correct answer, which is privileged information that does not exist at test time. This does not require a more capable model: we use the same model to generate solutions and synthetic rationales in the training data.
@@
Table A.3: Zero-shot prompt for our LLM-as-a-Judge evaluation results based on Gemini 1.0 Pro.
Prompt for LLM-as-a-Judge on GSM and MATH
You are a math teacher. Grade the Solution, verifying correctness step by step.
Question: {problem}
Solution: {solution}
@@ Appendix B Hyper-parameters for Verifier Training
For Gemma-based verifiers, we pick the best checkpoint based on validation accuracy of verification on held out problems and solutions. We always use data balancing between 50% correct solutions and 50% incorrect solutions in training.
@@
After doing a sweep of learning rates (LR), we find that an LR of $[2e-6,1e-6,5e-7]$ works well for our tasks considered (with LR=$2e-6$ generally being the best). We use a weight decay of $21e-$ 2, and do not apply any dropout. We use the Adam optimizer (Kingma, 2014) with decoupled weight decay (Loshchilov and Hutter, 2017) and a gradient norm clipping of $1.0$. We use a linear warmup of $1000$ gradient steps, and a cosine decay schedule that decays to $10\%$ of the peak learning rate after a decay period. We finetune for 300K steps with a batch size of 64, and use seqio (Roberts et al., 2022) library to create data mixtures.
@@
We finetune Gemma-based discriminative RMs by using a special token’s logit for classification. We chose the best performing ORM on our validation sets by launching a large sweep over learning rates $[1e-7,5e-7,1e-6,2e-6,3e-6,5e-6]$, weight decay $[1e-3,1e-2,1e-1]$ and dropouts $[1e-3,5e-3,1e-2,0]$. We also schedule the learning rate with a linear ramp up and a cosine decay. We use a Z-loss $=10^{-4}\cdot\log^{2}Z$ (where $Z$ is the softmax normalizer of all logits) for regularization purposes (Chowdhery et al., 2023; Wortsman et al., 2023). Results obtained with learning rate $71e-$ 7 and dropout=00.
DPO
We first finetune Gemma-based generative models using SFT on correct solutions to obtain a reference policy $\pi_{\text{ref}}$, and then initialize from this reference policy to train generator $\pi_{\text{DPO}}$ with the DPO loss on a dataset of pairs of correct and incorrect solutions. We conduct a hyper-parameter sweep for both the learning rate (LR) and the $\beta$ coefficient in DPO loss: for LR we sweeped $[1e-7,5e-7,1e-6,2e-6]$ and found $61e-$ 6 to work best; for $\beta$ we considered $[0.01,0.1,0.5,1.0,2.0]$ and used $0.1$. After DPO is trained, instead of using $r=\log\pi_{\text{DPO}}(\text{solution}\mid\text{question})-\log\pi_{\text{ref}}(\text{solution}\mid\text{question})$ as the score (as defined in DPO’s derivation), we find that directly the sequence log probability of the final DPO policy $\log\pi_{\text{DPO}}(\text{solution}\mid\text{question})$ as the score (without subtracting the log prob from reference policy) results in better performance in verification (see Figure D.5); a similar finding was also noted in (Hosseini et al., 2024).
Appendix C Additional Details
Data filtering for synthetic verification CoT
Since the answer checker (either based on string matching or Sympy library (Meurer et al., 2017)) is not perfect, there will inevitably be false negatives in the model-generated solutions. Besides, it is possible for a solution to arrive at the right answer with an incorrect reasoning path, so there will also be false positives in solutions. We use the following strategy to mitigate the issue of false negatives and false positives: when selecting the synthetic verification rationales (generated under reference guidance) for training, we only keep the rationales from solutions where more than 50% of verification rationales agree with the correctness returned by the answer checker.
Weighted Self-Consistency
typically sums the verifier scores (across solutions) for each answer, and picks the answer with the highest summed scores. We find that summing the top-K scores (rather than summing all scores) for each answer slightly improves performance. This means that for each answer, we only consider the correctness of its top-K solutions. We use K=6 for GSM and K=4 for MATH.
Appendix D Additional Results
Ablating generation loss weight ($\lambda$) in GenRM. Adding too much generation data negatively impacts verification, while intermediate values yield the best results, as shown in Figure D.3. By default, all GenRM experiments use unified training for verification with solution generation (5), with $\lambda=1/3$ for algorithmic tasks and $\lambda=1/4$ for GSM8K.
Data scaling for CoT verifiers. GenRM-CoT shows that the GenRM-CoT performance improves as we increase the number of solutions per problem from 8 to 32, in terms of RM accuracy and Best-of-N Accuracy, as shown in Figure D.3.
Figure D.1: GenRM (without using CoT) performs slightly better or comparable to Discriminative RM across different tasks, while outperforming DPO verifiers.
Figure D.2: Data scaling for GenRM-CoT on GSM8K with Gemma-7B. We observe that both the RM accuracy and Best-of-N performance improve as we scale up the number of rationales per solution and solutions per problem. When adding more solutions, we use 4 rationales per solution. Here, we compute GenRM-CoT scores with CoT rationales generated using greedy decoding, as discussed in (6).
Figure D.3: Impact of generation loss coefficient ($\lambda$) on GenRM verifier with Gemma-7B on GSM8K test results. Adding a solution generation loss ($\lambda>0$) can further help GenRM, with $\lambda=1/4$ being a good value for GSM.
Figure D.4: Weighted Self-Consistency on GSM8K: Unlike MATH, GSM8K shows no visible gain from Weighted SC: the percentage of problems solved increases only slightly from 93.4% to 93.5% with 16 solutions, likely because improvement potential has saturated.
Figure D.5: Ablation of DPO reward function. We find that directly the sequence log probability of the final DPO policy as the score (without subtracting the log prob from reference policy) results in better performance.
Appendix E Examples Verification rationales from GenRM-CoT: GSM8K Test and MATH500
Example of GenRM-CoT on Grade School Math
Discriminative RM score: 0.8414
GenRM-CoT: Let’s verify step by step.
**Verification:**
Step 1: Half of 4 hours is 2 hours.
Expected Answer: Correct.
Verification: Is the answer correct (Yes/No)? No
Gen-RM (Majority Voting) score: 0.2498
Table E.1: GenRM CoT Example 1
Example of GenRM-CoT on Grade School Math
@@ Appendix E Examples Verification rationales from GenRM-CoT: GSM8K Test and MATH500
Table E.2: GenRM CoT Example 2
@@
Table E.3: GenRM CoT Example 3
@@
Table E.4: GenRM CoT Example 4
@@
Table E.5: GenRM CoT Example 4 (Continued)
@@
Table E.6: GenRM CoT Example 5
@@
Table E.7: GenRM CoT Example 6
@@
Table E.8: GenRM CoT Example 7
@@
Table E.9: GenRM CoT Example 8
@@
Table E.10: GenRM CoT Example 9
@@
Table E.11: GenRM CoT Example 10
@@
Table E.12: GenRM CoT Example 11